OpenStudy (anonymous):

Circle A and circle B are congruent. CD is a chord of both circles If CD = 16 cm and the radius is 17 cm, how long is AB?

OpenStudy (superdavesuper):

Note that AB n CD bi-sect each other....let's call the point where they cross each other X....so AX=XB and CX=?

OpenStudy (anonymous):

DX?

OpenStudy (superdavesuper):

yup good job! as CD is given, u know CX=DX=?

OpenStudy (anonymous):

Got lost sorry CD=16 ...CX=DX= lost sorry

OpenStudy (superdavesuper):

look at the pic....CD=___ + ___ right?

OpenStudy (anonymous):

CD=16 +__ not sure what the next blank is

OpenStudy (superdavesuper):

Since we call the intersection point X, now the line CD is divided into two parts; CX and DX right? And u just said CX=DX....

OpenStudy (anonymous):

So I divided 16?

OpenStudy (superdavesuper):

yes divided it by 2 n u get CX

OpenStudy (anonymous):

8 ?

OpenStudy (superdavesuper):

hahaha why the ? com'on u KNOW 16/2=8 :P

OpenStudy (anonymous):

Lol just making sure I'm correct lol

OpenStudy (superdavesuper):

OK now look at the triangle AXC....notice <X is a right-angle....

OpenStudy (anonymous):

Yes I see that

OpenStudy (anonymous):

Yes I see that?

OpenStudy (superdavesuper):

So u know this property about right-angle triangle: a^2 + b^2 = c^2? where c is the longest side?

OpenStudy (anonymous):

Yes Pythagorean Therom familiar with it

OpenStudy (superdavesuper):

Cool so applying it for triangle AXC then....

OpenStudy (anonymous):

so am I doing 8^2 * 8^2 =c^2

OpenStudy (superdavesuper):

Nope...just write it out in terms of AX^2, CX^2 and AC^2 first plz?

OpenStudy (anonymous):

Okay I wrote down on my work out paper AX^2 * CX^2 AC^2

OpenStudy (superdavesuper):

U already know CX=8 and AC=raidus so look it up from ur Question...

OpenStudy (anonymous):

Cx=8 Ac=17

OpenStudy (superdavesuper):

so put them into ur eqn u get....

OpenStudy (anonymous):

18.7

OpenStudy (superdavesuper):

Nope set up ur eqn first plz?

OpenStudy (anonymous):

8^2 +17^2=

OpenStudy (superdavesuper):

Which side of the = sign AC^2 should be on? Left or right?

OpenStudy (anonymous):

Left?

OpenStudy (superdavesuper):

read Pythagorean Therom: AC is the longest side...

OpenStudy (anonymous):

Ohhh ... Right then !

OpenStudy (superdavesuper):

Right on :) soooooo.....

OpenStudy (anonymous):

Have to work backwards ......8^2+ ___=17^2

OpenStudy (superdavesuper):

yup yup

OpenStudy (anonymous):

15?

OpenStudy (superdavesuper):

that is the right answer :)

OpenStudy (anonymous):

thankyou ! so much

OpenStudy (superdavesuper):

welcome n thanks 4 working w me hehee